Find out whether or not there is a dependency relation between sales and store location for the following 50 stores (5% significance level). Sales/Location:High: Near Movie House is 5,Near House is 7, Near School is 6. Sales/Location:Average: Near Movie House is 7,Near House is 2, Near School is 6. Sales/Location:Low: Near Movie House is 1,Near House is 9, Near School is 7.


1
Expert's answer
2021-11-16T17:46:45-0500

Chi-Square test


"H_0:" sales and store location are independent

"H_a:" sales and store location are dependent


using online calculator www.socscistatistics.com, we get:


"\chi^2=\sum\frac{(O_i-E_i)^2}{E_i}"


where Oi is observed value,

Ei is expected of cell:


"E_i=\frac{(row\ total)(column\ total)}{total\ number}"


"df=(rows-1)(columns-1)"


The chi-square statistic is 8.8341. The p-value is .065383.

Since p-value > 0.05, we accept the null hypothesis.

Sales and store location are independent.